Performance Products Corporation makes two products, titanium Rims and Posts. Data regarding the two products follow:

   

Direct
Labor-Hours
per unit
Annual
Production
Rims 0.70 20,000 units
Posts 0.60 84,000 units

Additional information about the company follows:

  1. Rims require $16 in direct materials per unit, and Posts require $12.
  2. The direct labor wage rate is $15 per hour.
  3. Rims are more complex to manufacture than Posts and they require special equipment.
  4. The ABC system has the following activity cost pools:
Estimated Activity
Activity Cost Pool Activity Measure Estimated
Overhead
Cost
Rims Posts Total
Machine setups Number of setups $ 28,560 80 130 210
Special processing Machine-hours $ 159,600 6,000 0 6,000
General factory Direct labor-hours $ 680,000 10,000 30,000 40,000

Required:

1. Compute the activity rate for each activity cost pool. (Round your answers to 2 decimal places.)

Activity Cost Pool Activity Rate
Machine setups per setup
Special processing per MH
General factory per DLH
Answer #1
Estimated Activity
Activity Cost Pool Activity Measure Estimated
Overhead
Cost (i)
Rims Posts Total activity (ii) Activity rate (i)/(ii)
Machine setups Number of setups $ 28,560 80 130 210 28,500/210 = 136 per setup
Special processing Machine-hours $ 159,600 6,000 0 6,000 159,600/6,000 = 26.6 per MH
General factory Direct labor-hours $ 680,000 10,000 30,000 40,000 680,000/40,000 = 17 Per DLH
Activity Cost Pool Activity Rate
Machine setups 136 per setup
Special processing 26.6 per MH
General factory 17 per DLH

kindly give a positive rating if you are satisfied with the solution. Please ask if you have any query related to the question, Thanks.